OpenStudy (anonymous):

Use basic identities to simplify the expression. cos^2 θ / sin^2 θ + csc θ sin θ A. csc2θ B. sec2θ C. 1 D. tan2θ

OpenStudy (anonymous):

cos^2 /sin^2 = tan^2 csc *sin = 1 tan^2 + 1 = sec^2 put (theta) after any trigs, hope this help

OpenStudy (anonymous):

is it A

OpenStudy (anonymous):

it is B but you type sec 2(theta). it is not that. it is sec^2(theta)